# CLIPin: A Non-contrastive Plug-in to CLIP for Multimodal Semantic Alignment

### v1

@@ Introduction
For natural domain, we train on COCO (Lin et al. 2014) (82.8K images, 414.1K captions) and MUGE [^1] https://tianchi.aliyun.com/muge (250.4K image-text pairs from e-commerce). Evaluation is conducted on five benchmarks: ① CIFAR-10 (Krizhevsky and Hinton 2009), ② CIFAR-100 (Krizhevsky and Hinton 2009), ③ SUN397 (Xiao et al. 2016), ④ PASCAL VOC2007222http://www.pascal-network.org/challenges/VOC/voc2007/workshop/index.html, and ⑤ Caltech-101 (Fei-Fei, Fergus, and Perona 2004). For medical domain, we train on a private dataset (Tongren) with 451.9K retinal image–report pairs from Beijing Tongren Hospital, and evaluate on ⑥ RFMiD (Pachade et al. 2021), ⑦ ODIR [^3] https://odir2019.grand-challenge.org, ⑧ REFUGE (Orlando et al. 2020), ⑨ MESSIDOR (Decencière et al. 2014), and ⑩ FIVES (Jin et al. 2022).
@@
All models adopt ViT-B/16 (Dosovitskiy et al. 2021) as the image encoder. Models trained on COCO are initialized with CLIP, while those on MUGE and Tongren use CN-CLIP (Yang et al. 2022). The text encoder varies across datasets but is fixed per experiment. Input images are resized to $224\times 224$, randomly horizontal flipped (probability $0.5$), and augmented with color jitter (strength $0.1$). The max text length $l$ is 77. We use AdamW (Loshchilov and Hutter 2018) with a learning rate of $3\times 10^{-5}$, warmup of $100$ iterations, $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1\times 10^{-6}$, and weight decay $\lambda=0.001$. The momentum coefficient $\beta=0.95$, temperature $\tau=0.07$, and weighting coefficients $\lambda_{\mathrm{inter}}$ and $\lambda_{\mathrm{intra}}$ are initialized to $1.0$. The batch size $B=256$. The training takes approximately 24 hours on a single RTX 3090 GPU using automatic mixed precision, with a memory consumption of 14 GB.
@@
Our method is evaluated using linear probing and prompt-based out-of-distribution zero-shot classification (prompt-based OOD-ZSC). Linear probing follows (He et al. 2022), training a linear classifier atop frozen encoders to assess representation quality. Prompt-based OOD-ZSC evaluates zero-shot transfer by computing image–text feature similarity, using category prompts as text labels. This evaluates both generalization and modality alignment. As all datasets are multi-labeled, we report Area Under the ROC Curve (AUC) and mean Average Precision (mAP), where AUC reflects global discriminative power and mAP captures performance on long-tailed labels. For qualitative analysis, we use multimodal Grad-CAM (Selvaraju et al. 2017) to generate heatmaps conditioned on text inputs.
@@
Comparative study
Linear probing classification.
We compare the linear probing performance of CLIP(Radford et al. 2021), xCLIP (Zhou et al. 2023), and CLIP intergated with CLIPin (Ours). CLIP serves as the baseline, while xCLIP represents a state-of-the-art fusion method that introduces a non-contrastive auxiliary loss to enhance contrastive learning. All models are trained from scratch under a unified setup. As shown in Table 1 (left), CLIPin consistently improves AUC and mAP across datasets, with notable gains in challenging categories.
Prompt-based OOD-ZSC classification.
Generalization study of CLIPin.
To evaluate the effectiveness and plug-and-play feasibility of the proposed CLIPin, we selected several state-of-the-art methods known for enhancing the robustness of contrastive learning: ALBEF (Li et al. 2021a), BLIP (Li et al. 2022), and CoCa (Yu et al. 2022). ALBEF improves vision-language pretraining via momentum-based feature alignment and contrastive objectives; BLIP leverages bootstrapped captions and weakened supervision signals to enrich visual-language alignment; CoCa combines contrastive and generative learning in a unified multimodal framework. All models are trained from scratch to ensure a fair comparison. We integrated CLIPin into their contrastive learning modules and compared the linear probing classification performance before and after this integration, as shown in Table 2, to demonstrate that CLIPin can further enhance these frameworks.
@@
Ablation study
To assess the contribution of each component in CLIPin, we perform ablation studies in Table 3 using the COCO and Tongren as training datasets, evaluating linear probing classification performance on two downstream benchmarks: PASCAL VOC2007 and RFMiD. Starting from a baseline CLIP model, we add CLIPin’s key modules: inter-modal alignment, intra-modal alignment, and pre-projector sharing sequentially, and analyze the impact of each.
The results reveal several noteworthy trends. First, incorporating inter-modal alignment alone provides marginal improvements and may even degrades the performance slightly, suggesting that isolated cross-modal alignment, especially when implemented via a momentum-based target encoder, may introduce instability in the early training stage. The lack of anchoring in the unimodal space makes it harder to form robust semantic correspondences across modalities. Introducing intra-modal alignment alleviates these issues, leading to clearer gains across tasks. Finally, adding the shared pre-projectors further boosts the performance, confirming that unifying parts of the architecture across learning paradigms does not interfere with, and may even synergize dual training objectives. This validates the effectiveness of CLIPin’s plug-in design, showing that its benefits arise not only from isolated modules but also their joint interaction.
Multimodal Grad-CAM visualization
@@
To illustrate how CLIPin enhances feature interpretability more intuitively, we adopt multimodal Grad-CAM for visualization. In natural domain, the model is trained on COCO and evaluated on PASCAL VOC2007; in medical domain, it is trained on Tongren and evaluated on FGADR (Zhou et al. 2020), which includes pixel-level lesion annotations to enable a precise assessment of whether the activated regions are correspond to the pathological areas. As shown in Fig. 2, we compare Grad-CAM maps generated from models trained with CLIP, xCLIP, and CLIP with CLIPin.
@@
In natural domain (column “Train”–“Bird”), CLIP with CLIPin yields denser and more spatially continuous activations that follow the shape and boundaries of target objects, while suppressing irrelevant background signals. In medical domain (column “Cotton wool spots”–“Hemorrhage”), CLIPin improves text-to-visual attention significantly, enabling more accurate localization of lesion areas in appearance, position, and spatial extent, with better correspondence to expert annotations. The improved localization and semantic focus suggest that CLIP with CLIPin captures domain-specific visual cues better, which is due to the instance-level supervision from the non-contrastive component. These qualitative results reinforce our quantitative findings: CLIPin not only boosts performance metrics but also enhances the interpretability, semantic consistency, and zero-shot generalization of the learned representations.
Conclusion
@@
We propose CLIPin, a unified non-contrastive plug-in that enhances multimodal semantic alignment and can be seamlessly integrated into existing contrastive learning pipelines, functioning as a plug-and-play module that improves representation quality, generalization and cross-modal alignment. By introducing non-contrastive pathways, CLIPin addresses the key limitations of CLIP-style models, such as semantic looseness and redundancy. Extensive experiments demonstrate that CLIPin outperforms prior methods and improves the performance across diverse architectures consistently. Although CLIPin has a cyclic and modality-symmetric design that can be naturally extended to more than two modalities, this work focuses on the image–text setting due to practical constraints. Future work will explore scaling to larger multimodal corpora and further investigating the synergy between contrastive and non-contrastive paradigms.

### v2

@@ 4.1 Experiment settings
For natural domain, we train on COCO [37] (82.8K images, 414.1K captions) and MUGE [^1] https://tianchi.aliyun.com/muge (250.4K image-text pairs from e-commerce). Evaluation is conducted on five benchmarks: ① CIFAR-10 [38], ② CIFAR-100 [38], ③ SUN397 [39], ④ PASCAL VOC2007222http://www.pascal-network.org/challenges/VOC/voc2007/workshop/index.html, and ⑤ Caltech-101 [40]. For medical domain, we train on a private dataset (Tongren) with 451.9K retinal image–report pairs from Beijing Tongren Hospital, and evaluate on ⑥ RFMiD [41], ⑦ ODIR [^3] https://odir2019.grand-challenge.org, ⑧ REFUGE [42], ⑨ MESSIDOR [43], and ⑩ FIVES [44].
@@
All models adopt ViT-B/16 [45] as the image encoder. Models trained on COCO are initialized with CLIP, while those on MUGE and Tongren use CN-CLIP [46]. The text encoder varies across datasets but is fixed per experiment. Input images are resized to $224\times 224$, randomly horizontally flipped (probability $0.5$), and augmented with color jitter (strength $0.1$). The max text length $l$ is 77. We use AdamW [47] with a learning rate of $3\times 10^{-5}$, warmup of $100$ iterations, $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1\times 10^{-6}$, and weight decay $\lambda=0.001$. The momentum coefficient $\beta=0.95$, temperature $\tau=0.07$, and weighting coefficients $\lambda_{\mathrm{inter}}$ and $\lambda_{\mathrm{intra}}$ are initialized to $1.0$. The batch size $B=256$. The training takes approximately 24 hours on a single RTX 3090 GPU using automatic mixed precision and gradient checkpointing, with a memory consumption of 14 GB.
@@
We evaluate using linear probing and prompt-based out-of-distribution zero-shot classification (prompt-based OOD-ZSC). Linear probing follows He et al. [48], training a linear classifier atop frozen encoders. Prompt-based OOD-ZSC evaluates zero-shot transfer by computing image-text feature similarity, using category prompts as text labels (translated into accurate Chinese terms when evaluating models trained in Chinese). This evaluates both generalization and modality alignment. As all datasets are multi-labeled, we report Area Under the ROC Curve (AUC) and mean Average Precision (mAP), where AUC reflects global discriminative capability and mAP captures performance on long-tailed labels. The reported values are averaged over five repeated runs. Throughout the experiments, we assess statistical significance using paired t-tests, all improvements reported are statistically significant with $p<0.05$. For qualitative analysis, we use multimodal Grad-CAM [49] to generate heatmaps conditioned on text inputs.
4.2 Comparative study
Linear probing classification.
We compare the linear probing performance of CLIP[1], xCLIP [10], and CLIP intergated with CLIPin (Ours). CLIP serves as the baseline, while xCLIP represents a state-of-the-art fusion method that introduces a non-contrastive auxiliary loss to enhance contrastive learning. All models are trained from scratch under a unified training setup. As shown in Table 1 (left), CLIPin consistently improves both AUC and mAP across datasets, with notable gains in challenging categories.
Prompt-based OOD-ZSC classification.
Generalization study of CLIPin.
To evaluate the effectiveness and plug-and-play feasibility of the proposed CLIPin, we selected several state-of-the-art methods known for enhancing the robustness of contrastive learning: ALBEF [11], BLIP [12], and CoCa [24]. ALBEF improves vision-language pretraining via momentum-based feature alignment and contrastive objectives; BLIP leverages bootstrapped captions and weakened supervision signals to enrich visual-language alignment; CoCa combines contrastive and generative learning in a unified multimodal framework. All models are trained from scratch to ensure a fair comparison. We integrated CLIPin into their contrastive learning modules and compared the linear probing classification performance before and after this integration, as shown in Table 2, to demonstrate that CLIPin can further enhance these advanced frameworks.
4.3 Ablation study
To assess the contribution of each component in CLIPin, we perform ablation studies in Table 3 using the COCO and Tongren as training datasets, evaluating linear probing classification performance on two downstream benchmarks: PASCAL VOC2007 and RFMiD. Starting from a baseline CLIP model, we incorporate the three key modules of CLIPin: inter-modal alignment, intra-modal alignment, and pre-projector sharing sequentially, and analyze the impact of each.
The results reveal several noteworthy trends. First, incorporating inter-modal alignment alone provides marginal improvements, and in some cases even slightly degrades the performance. This suggests that isolated cross-modal alignment, especially when implemented via a momentum-based target encoder, may introduce instability or convergence difficulties in the early training stage. The lack of anchoring in the unimodal space makes it harder for the model to form robust semantic correspondences across modalities. Introducing intra-modal alignment alleviates these issues, leading to clearer gains across tasks. Finally, adding the two shared pre-projectors further boosts the performance, confirming that unifying parts of the architecture across learning paradigms does not interfere with, and may even synergize dual training objectives. This validates the effectiveness of the plug-in design of CLIPin, demonstrating that its benefits not only rely on isolated modules but also emerge from their joint interaction.
@@ 4.3 Ablation study
4.4 Multimodal Grad-CAM visualization
@@
To illustrate how CLIPin enhances the interpretability of features learned by CLIP more intuitively, we adopt multimodal Grad-CAM for visualization. In natural domain, the model is trained on COCO and evaluated on PASCAL VOC2007. In medical domain, the model is trained on Tongren and evaluated on FGADR [50], which includes pixel-level lesion annotations to enable a precise assessment of whether the activated regions are correspond to the pathological areas. As shown in Fig. 2, we compare Grad-CAM activation maps generated from models trained with CLIP, xCLIP, and CLIP with CLIPin.
@@
In natural domain (column “Train”–“Bird”), CLIP with CLIPin yields denser and more spatially continuous activations that follow the shape and boundaries of target objects, while suppressing irrelevant background signals. In medical domain (column “Cotton wool spots”–“Hemorrhage”), CLIPin improves the alignment of text to visual attention significantly, enabling more accurate localization of lesion areas in terms of appearance, position, and spatial extent, with better correspondence to expert-annotated ground truth. The improved localization and semantic focus suggest that CLIP with CLIPin is better equipped to capture domain-specific visual cues, which is due to the additional instance-level supervision introduced by the non-contrastive component. These qualitative results further support our quantitative findings: integrating CLIPin into CLIP not only boosts performance metrics but also enhances the interpretability, semantic consistency, and zero-shot generalization of the learned representations.
5 Conclusion
@@
In this work, we propose CLIPin, a unified non-contrastive plug-in designed to enhance multimodal semantic alignment. CLIPin can be seamlessly integrated into existing contrastive learning pipelines, functioning as a plug-and-play module that improves representation quality, generalization and cross-modal alignment. By introducing additional non-contrastive pathways, CLIPin addresses the key limitations of CLIP-style models, such as semantic looseness and redundancy. Extensive experiments demonstrate that CLIPin outperforms prior methods and delivers robust performance gains across diverse architectures consistently. Although CLIPin is implemented with a cyclic and modality-symmetric design that can be naturally extended to more than two modalities, this work focuses on the image–text setting due to practical constraints. Extending CLIPin beyond the bimodal case remains a promising direction for future research. We also plan to further explore the synergy between contrastive and non-contrastive paradigms, including improving robustness to data augmentations and scaling to larger multimodal corpora.
Appendix A Appendix
A.1 Data collection and curation
In this work, we position our method within the general medical domain. For empirical validation, we focus on ophthalmology, where we constructed a large-scale dataset of retinal images paired with diagnostic reports. The dataset Tongren was collected through a large-scale telemedicine initiative involving 172 hospitals across mainland China between 2012 and 2020. Patient submissions included demographic information, clinical complaints, medical histories, ophthalmic measurements (e.g., visual acuity and intraocular pressure), and retinal images. Experienced ophthalmologists subsequently reviewed these cases and provided diagnostic reports.
To the best of our knowledge, no comparable public resource of retinal image-text pairs currently exists. Due to strict privacy regulations and institutional policies, the dataset cannot be publicly released, despite thorough de-identification. While our approach is modality-agnostic and can readily extend to public datasets in other medical domains, this paper focuses on the ophthalmology case study, and we consider broader multi-domain validation an important direction for future work.
Data acquisition.
In total, we obtained over 900K ophthalmic images from more than 400K patient visits. A multi-stage curation pipeline was implemented to ensure data quality and consistency, combining automated classification, report-based filtering, and redundancy reduction. After this process, we retained approximately 451.9K high-quality retinal images, each paired with a corresponding diagnostic text.
Report structure.
Each diagnostic report contains two major components: (i) findings, which describe observable fundus features and abnormalities; and (ii) impression, summarizing the case into a preliminary diagnosis and clinical recommendation.
This pipeline yielded a modality-consistent dataset suitable for model training and evaluation, while ensuring compliance with both ethical and privacy requirements.
A.2 Training dynamics of inter- and intra-modal alignment weights
Figure 3: Training dynamics of $\lambda_{\mathrm{inter}}$ and $\lambda_{\mathrm{intra}}$.
We collect the values of weights $\lambda_{\mathrm{inter}}$ and $\lambda_{\mathrm{intra}}$ throughout training and plot their trajectories over training steps, as shown in Fig. 3. The curves cover the training period up to the point where the model reaches peak evaluation performance, after which overfitting begins and the dynamics of both weights remain stable. It can be observed that $\lambda_{\mathrm{inter}}$ increases steadily, while $\lambda_{\mathrm{intra}}$ remains relatively stable. This suggests that, as training progresses, inter-modal alignment gradually becomes the primary driver of representation learning, surpassing contrastive learning and intra-modal alignment in contributing to performance gains.
A.3 Additional ablation studies
Ablation of output dimensionality of the shared pre-projectors.
Table 4: Ablation study on the output dimensionality of shared pre-projectors
Output dimension of pre-projectors 512 1024 2048
AUC/mAP, % 87.22/43.16 87.43/43.43 87.22/41.84
Table 5: Ablation study on varying batch sizes $B$ (AUC/mAP, %)
$B$ CLIP Ours
128 86.65/41.36 87.49/43.33
256 87.18/41.92 87.43/43.43
512 87.08/42.34 87.89/43.70
To evaluate the impact of the output dimensionality of the shared pre-projectors, we conduct an ablation study by varying the output dimensions across a range of values (512, 1024, and 2048) while keeping all other hyperparameters fixed. The models are trained on COCO [37] dataset and evaluated on PASCAL VOC2007 using linear probing classification. As summarized in Table 4, we observe that the configuration used in our main experiments corresponds to the optimal output dimensionality. Increasing or decreasing the dimension beyond this setting leads to performance degradation, suggesting that our chosen configuration strikes a good balance between representation quality and optimization stability.
Ablation of varying batch sizes.
To investigate whether integrating CLIPin into CLIP [1] reduces the reliance on large batch sizes, we compare the performance of the original CLIP and CLIP with CLIPin under varying batch sizes $B$ of 128, 256, and 512. We use COCO as the training dataset and evaluate the linear probing performance on PASCAL VOC2007. Results are reported in Table 5. We observe that across different batch size settings, CLIPin consistently improves the quality of the learned representations.
Table 6: Ablation study on image and text augmentation (AUC/mAP, %)
Image augmentation
Text augmentation
Linear probing 87.42/42.67 87.43/43.43 87.48/42.71 87.63/43.22
Prompt-based OOD-ZSC 79.50/29.03 82.13/32.20 81.92/30.74 81.01/31.98
Table 7: Ablation study on inter- and intra-modal alignments
Inter-modal Alignment CLIP with CLIPin
Intra-modal Alignment
AUC/mAP, % 85.43/35.85 75.08/21.22 87.42/41.79 87.43/43.43
Ablation of data augmentation.
To evaluate the impact of data augmentation on model performance, we disable either image augmentation (i.e., setting $I^{\mathrm{(1)}}$ = $I^{\mathrm{(2)}}$) or text augmentation (i.e., setting $\hat{T}^{\mathrm{(1)}}$ = $\hat{T}^{\mathrm{(2)}}$) selectively, and compare the results against the baseline where both image and text augmentation are disabled. Given that the COCO dataset provides rich textual augmentation (i.e., multiple captions per image), we use it as the training dataset and evaluate linear probing and prompt-based OOD-ZSC classification performance on PASCAL VOC2007. Results are shown in Table 6. We observe that enabling image augmentation improves both visual representation learning and multimodal semantic alignment substantially. In contrast, the benefits of text augmentation are relatively limited and, in some cases, may even diminish the gains introduced by image augmentation. Therefore, in our main experiments, we only enable image augmentation to maximize overall performance.
Ablation of modal alignments.
To further investigate the respective contributions of non-contrastive inter- and intra-modal alignment, we train CLIPin independently without CLIP loss and conduct ablation studies by disabling each alignment component selectively. Specifically, we compare three variants: using only inter-modal alignment, using only intra-modal alignment, and using both. We use COCO as the training dataset and evaluate the linear probing performance on PASCAL VOC2007. The ablation results of CLIPin with different alignment configurations are reported in Table 7, with CLIP with CLIPin included as a reference. The results reveal that using only inter-modal alignment leads to poor representation quality, consistent with our earlier analysis that multimodal heterogeneous encoders are prone to collapse when optimized without intra-modal regularization. On the other hand, using only intra-modal alignment also results in poor performance due to the absence of multimodal supervision, which makes the encoders overly sensitive to data augmentations. Combining intra- and inter-modal alignment mitigates these issues significantly, yielding results that are close to those obtained with contrastive learning.
Combining insights from Fig. 3 and Table 7, we conclude that CLIPin and contrastive learning complement each other. In the early stages of training, the clear optimization signals from contrastive learning and intra-modal alignment help stabilize and guide the inter-modal non-contrastive objective. As training progresses, inter-modal alignment gradually takes over as the primary driver of representation learning and semantic alignment, compensating for the limitations of contrastive modeling effectively.
A.4 Additional generalization studies of CLIPin
Table 8: Generalization study of CLIPin on downstream benchmarks (AUC/mAP, %)
Linear probing
ALBEF BLIP CoCa OTTER SLIP
88.02/43.52 88.51/46.06 86.06/38.85 81.94/30.62 87.78/43.18
+CLIPin 88.14/44.99 88.95/47.68 86.03/39.77 82.23/31.38 88.10/43.88
Prompt-based OOD-ZSC
ALBEF BLIP CoCa OTTER SLIP
83.85/33.95 82.88/35.58 78.59/27.12 77.27/22.89 80.93/27.77
+CLIPin 83.89/36.05 83.94/35.80 78.70/27.18 78.15/23.67 80.40/28.04
To further support the generalization capability of CLIPin, we build upon our main experimental setup and integrate CLIPin into five representative vision-language training frameworks that incorporate contrastive learning: ALBEF [11], BLIP [12], CoCa [24], OTTER [13], and SLIP [51]. All models are trained from scratch on COCO and evaluated on PASCAL VOC2007. For each model, we compare the linear probing and prompt-based OOD-ZSC performance with and without CLIPin. Detailed results are presented in Table 8. The results further demonstrate the generalization ability of CLIPin and show that integrating CLIPin consistently enhances multimodal semantic alignment capabilities. These findings reinforce our conclusion that CLIPin mitigates the key limitations (semantic looseness and redundancy) of the CLIP framework effectively, while offering clear advantages in zero-shot multimodal semantic alignment under distribution shift.
A.5 Detailed results by category on downstream benchmarks
To demonstrate the improvement in representation learning quality brought by CLIPin more concretely, we compare CLIP, xCLIP [10], and CLIP with CLIPin (Ours) on a per-category evaluation using both linear probing and prompt-based OOD-ZSC. The models are trained on COCO and evaluated on PASCAL VOC2007. Evaluation metrics include AUC and Average Precision (AP). The detailed results are illustrated in Table 9. Categories A-T correspond to “aeroplane”, “bicycle”, “bird”, “boat”, “bottle”, “bus”, “car”, “cat”, “chair”, “cow”, “diningtable”, “dog”, “horse”, “motorbike”, “person”, “pottedplant”, “sheep”, “sofa”, “train”, and “tvmonitor”. The results show that the performance gains introduced by CLIPin are consistent across the vast majority of categories, rather than being concentrated in a few outliers. Moreover, CLIP with CLIPin generally outperforms xCLIP. These findings provide further evidence of the robustness and effectiveness of CLIPin in enhancing representation quality.
Table 9: Per-category results on downstream benchmarks (AUC/AP, %)
Linear probing Prompt-based OOD-ZSC
CLIP xCLIP Ours CLIP xCLIP Ours
A 93.49/61.19 92.86/61.23 94.08/67.19 82.56/11.66 85.84/29.48 87.89/18.56
B 86.54/33.80 82.53/30.37 85.79/33.68 82.12/28.06 75.87/15.84 83.38/25.91
C 84.46/33.79 83.70/33.11 85.17/33.87 82.43/27.92 81.29/30.89 82.08/31.36
D 91.00/51.41 88.94/53.47 92.13/53.16 90.15/37.74 90.49/46.52 91.03/40.48
E 80.20/15.63 80.40/15.68 78.18/13.39 61.29/6.35 57.54/6.73 74.40/9.35
F 87.96/31.10 88.59/28.21 89.24/32.71 85.17/33.75 84.49/32.14 87.85/35.65
G 86.38/56.51 85.47/55.29 86.56/57.40 83.86/49.94 81.58/43.96 84.76/55.05
H 86.95/36.75 85.74/35.75 86.36/35.62 83.80/33.27 84.52/31.93 85.19/33.78
I 84.32/46.16 84.34/44.66 85.73/46.33 79.65/29.46 82.88/40.73 83.52/44.19
J 88.84/22.65 84.97/20.47 89.09/23.57 84.72/14.00 82.65/15.40 88.53/21.52
K 89.89/34.57 88.33/31.71 90.59/39.24 77.87/21.41 81.14/14.79 85.03/37.00
L 80.14/31.12 81.30/30.10 82.93/32.49 78.47/29.48 78.88/29.33 81.27/30.85
M 91.95/61.96 90.42/58.84 91.81/61.45 87.29/35.92 88.90/42.10 88.42/36.70
N 86.62/40.33 85.34/37.41 87.19/42.52 81.15/24.46 78.45/23.30 87.35/36.18
O 83.55/81.56 82.74/81.52 83.61/82.29 73.07/71.56 72.49/69.62 73.51/72.33
P 83.53/28.95 80.41/25.49 84.54/29.50 71.82/15.38 59.06/10.76 51.91/9.33
Q 89.15/40.98 90.18/31.91 89.56/39.37 89.69/30.39 88.37/27.30 87.52/30.70
R 88.00/36.98 85.37/28.01 88.14/36.52 83.22/33.24 83.44/28.98 86.74/31.77
S 92.87/52.36 91.95/50.78 93.45/57.55 88.40/38.01 87.48/38.21 87.89/46.98
T 87.19/39.25 87.55/36.98 88.76/42.63 71.28/15.61 75.82/16.77 86.41/39.59